OpenStudy (anonymous):

how to solve a cubic equation? Please explain in a simple manner because I have trouble in understanding them and what is this Synthetic division????? How is it helpful??

OpenStudy (lgbasallote):

here's a general concept..find a root then use synthetic division to divide the given polynomial by that root...the quotient will give you a quadratic equation,.you can use quadratic formula to find the remaining two roots

OpenStudy (lgbasallote):

now...going more specifically..synthetic division is a way of dividing polynomials

OpenStudy (ghazi):

to solve a cubic equation there are lot of methods... here you can find details http://en.wikipedia.org/wiki/Cubic_function

OpenStudy (anonymous):

wouldn't quadratic formula be easy

OpenStudy (ghazi):

well quadratic only suits when it can be factored easily...like in case X^3-7X+9=0 we can't apply quadratic

OpenStudy (anonymous):

Please explain step by step the process of finding the root of the cubic polynomial which of the easiest way

OpenStudy (anonymous):

Actually I would say..... to use FACTOR theorem and Remainder theorem is better

OpenStudy (lgbasallote):

actually...remainder/factor theorem + rational roots theorem...are you familiar with either terms?

OpenStudy (anonymous):

I remember using them but its been a long gap so please explain again

OpenStudy (anonymous):

the truth is that there is no easy way

OpenStudy (anonymous):

if the polynomial has been cooked up to make it easy, that is, if it has been specifically constructed to have integer or rational roots, then you can find them using the rational root theorem. but if you just write down some random third degree polynomial, you will have to use a long and ugly formula to find the roots. i believe it is in the wiki link above

OpenStudy (amistre64):

one way, if given the form: x^3 + ax^2 + bx + c = n is to eliminate the x^2 by a suitable replacement of: x = y-a/3 this will get you the form: y^3 + py = q; which somehow is simpler to deal with :/

mathslover (mathslover):

I will prefer you the way of : factor theorem

mathslover (mathslover):

\[\large{x^3+ax^2+bx+c}\] find the factors of c :\(\large{\textbf{plus and minus factors are included}}\) then by common sense and trial and error find only 1 zero of the given polynomial let us suppose we have the polynomial p(x) and a zero of this polynomial (a) so (x-a) will be the factor of p(x) (x-a)(b)=p(x) where b is another factor of p(x) , b may include 2 factors or more.. divide p(x) by (x-a) by long division or any other method you will get a polynomial suppose it is (x-b) hence we have (x-a)(x-b) =p(x) that is (x-a) and (x-b) are the factors of p(x)

OpenStudy (amistre64):

as satellite has already pointed out, that will only work if the poly has been constructed to conform to that method. rational roots test is effective only if there exists rational roots to the poly.

OpenStudy (amistre64):

you can do trial and error tests with consecutive integers (a < b) such that: f(a) < 0 and f(b)>0 this at least lets you know a root exists somewhere in the interval (a,b)

mathslover (mathslover):

That is a good point though :)

OpenStudy (anonymous):

can you guys teach me this with an example please..it would be helpful for me very much..please

OpenStudy (anonymous):

see you can find the first root by hit and trial and rest two roots by horners method or factorisation

OpenStudy (anonymous):

eg: \[x ^{3}-7x+6=0\]

OpenStudy (anonymous):

are you there?????

OpenStudy (anonymous):

what next @nitz

OpenStudy (anonymous):

so first root of this equation is 1 as (1)^3-7(1)+6=-6+6=0 thus it satisfies the equation ...... so by hit and trial .the first root is 1 right?????

OpenStudy (anonymous):

basically in hit and trial,we substitute values -1,0,1,.. and check out which satisfies the equation....generally you need to check just -1,1,2,3,4.....just starting nos.

OpenStudy (anonymous):

ok got it then

OpenStudy (anonymous):

no actually i was trying to figure out how to show you synthetic division here.........

OpenStudy (anonymous):

hmm can you please draw it

OpenStudy (anonymous):

can you please check this link...here is another example

OpenStudy (anonymous):

its not possible to draw...if u have any problem i can tell sorry

mathslover (mathslover):

what to draw i can..

OpenStudy (anonymous):

synthetic division method...

mathslover (mathslover):

what is the dividend and divisor @nitz ?

OpenStudy (anonymous):

the equation is above x^3-7x+6=0

OpenStudy (anonymous):

one root is 1

mathslover (mathslover):

oh k so this means x-1 is a factor of x^3-7x+6

OpenStudy (anonymous):

got it guys thanks @nitz